OpenStudy (anonymous):

If 3a + 2b = 7 and 2a + 2b = 9 , what is the value of 5a+4b/2

OpenStudy (raden):

3a + 2b = 7 2a + 2b = 9 add both equation above, get 5a + 4b = 16 then divided by 2 both side, get (5a+4b)/2 = 16/2 = ...... ?